### v1

# Evidence of equilibrium dynamics in human social networks evolving in time

Abstract
The dynamics of personal relationships remain largely unexplored due to the inherent difficulties of the longitudinal data collection process. In this paper, we analyse a dataset tracking the temporal evolution of a network of personal relationships among 900 people over the course of four years. We search for evidence that the network is in equilibrium, meaning that all macroscopic properties remain constant, fluctuating around stable values, while the internal microscopic dynamics are active. We find that the probabilities governing the network dynamics are stationary over time and that the degree distributions, as well as edge and triangle abundances match the theoretical equilibrium distributions expected under these dynamics. Furthermore, we verify that the system satisfies the detailed balance condition, with only minor point deviations, confirming that it is indeed in equilibrium. Remarkably, this equilibrium persists despite a high turnover in network composition, suggesting that it is an inherent characteristic of human social interactions rather than a trait of the individuals themselves. We argue that this equilibrium may be a general feature of human social networks arising from the competition between different dynamical mechanisms and also from the cognitive and material resources management of individuals. From a practical perspective, the fact that networks are in equilibrium could simplify data collection processes, validate the use of cross-sectional data-based methods like Exponential Random Graph Models, and inform the design of interventions. Our findings advance the understanding of collective human behaviour predictability and our ability to describe it using simple mathematical models.
Keywords Personal Relationships  $\cdot$ Social Network Analysis  $\cdot$ Longitudinal Data  $\cdot$ Equilibrium  $\cdot$ Network Dynamics
1 Introduction
Human social behaviour is driven by a complex interplay of cognitive, emotional, and social factors, which together shape the structure and dynamics of social networks. Focusing on personal relationships, prior research has shown that individuals tend to maintain a finite number of social bonds constrained by their cognitive capacity [1], and how these social bonds are structured is explained by the way humans allocate their cognitive resources to create functional relationships able to cover their social necessities [2, 3]. From a dynamical point of view, individuals continually adjust their social ties in response to emotional, environmental, and cognitive changes. Still, this dynamic behaviour of social bonds is less understood, mainly due to a lack of rich longitudinal data: longitudinal studies on social networks evolution are rare and often face issues like small sample sizes, noisy data, and limited conclusions. Our work focuses on the temporal evolution of social relationships and addresses the question of whether these complex tendencies give rise to stable trends.
Due to the interconnected nature of social ties, network theory constitutes an appropriate framework to investigate such phenomena. In network science [4], and in particular in the context of social networks [5], personal relationship networks are defined as networks where nodes represent people and links represent personal relationships, such as friendships or enmities. In general, data on these networks is gathered by selecting a sample population of unrelated individuals and asking each person to name the people with whom they have a relationship [6] and how they are related, or by surveying individuals within a closed environment, such as a school, university, or company [7]. Examples of datasets collected using the first strategy include the Caen Panel Survey [8], analysed extensively in [9, 10, 11, 12], and other similar datasets replicated on a smaller scale, like [13]. Regarding the second strategy, important datasets include Newcomb’s Fraternity data [14, 15, 16, 17, 18], and Sampson’s data [19]. Several studies have also collected data in school environments [20, 21, 22, 23, 24, 25], being specially relevant the National Longitudinal Study of Adolescent to Adult Health (Add Health) [26, 27]. These datasets represent an outstanding effort to overcome the challenges of collecting longitudinal data. However, they have limitations that constrain the conclusions that can be drawn from their analysis, aside from small sample sizes (with the exceptions of [24, 27]). The first method’s data are usually undirectional and binary, as the respondent reports the mere existence of third-party relationships according to their perception. The second method provides directed but still binary and partial data, often limited to within-class relationships with an artificial cut-off (of typically 5) to the number of relationships reported by the respondent imposed by the surveyor. Furthermore, many studies collected only two to three snapshots, limiting long-term dynamic analysis.
Given these limitations, some alternative ways to investigate the temporal evolution of personal relationships have been attempted. A possible strategy involves inferring the network of personal relationships from indirect data, such as phone calls [28], email exchanges [29] or face-to-face interaction data [30, 31], instead of using surveys. The only drawback is that conclusions about personal networks depend on the assumption that this indirect data is indeed a proxy of the underlying personal relationships. Other researchers have chosen to study network dynamics indirectly using cross-sectional data, which is easier to obtain. The basic idea behind this approach is that the dynamical mechanisms driving the temporal evolution of a network leave a fingerprint in the structure observed at single points in time. For instance, if nodes tend to reciprocate links, a static network snapshot will show a higher number of reciprocal links than expected by chance. This is the key idea behind Exponential Random Graph Models [32], which infer network formation mechanisms based on the observed structure, assuming that the observed network is a random sample from a family of networks produced by the model, and behind other similar methods [33]. Finally, a third strand of research has been more focused on developing models with specific hypothetical mechanisms and simulating the network’s evolution according to these models [34, 35, 36]. Closely related, some works make use of traditional statistical mechanics of the equilibrium techniques to explore structural patterns in networks [37, 38].
Using these data and methods, several studies have focused on specific aspects of the evolution of social ties. For instance, Rivera et al. explore the temporal dynamics of dyadic relationships [39], while Yap and Harrigan assess how various mechanisms can explain the observed behaviour across different datasets [40]. In contrast, our work takes a broader approach, focusing on the collective behaviour of social relationships. Specifically, we analyse whether personal relationship networks can exhibit stable trends over time, a phenomenon we refer to as equilibrium dynamics. In the following sections, we will provide a rigourous definition of this concept inherited from statistical physics. In short, we consider a social network to be in equilibrium when its macroscopic, average properties – such as the distribution of personal contacts; the prevalence of patterns like edges, triangles, or larger structures; and aggregate metrics like network density – remain constant, fluctuating around stable values, even as individual links continuously evolve at a microscopic level. Our interest is driven by the lack of a rigourous exploration within the framework of equilibrium dynamics of some of the mentioned studies [20, 23, 17, 25, 29, 28, 13, 41]. In particular, and interestingly, there are apparent trends of stability in other aspects of human social behaviour, which suggest a natural tendency toward the formation of stable patterns and self-reinforcing dynamics. For example, Hobbs and Burke show how social connections recover after the death of a friend [42], while Alessandreti et al. found that certain mobility patterns are conserved over time [43], linking these patterns to individuals’ social ties. Other studies, such as [44, 45, 46], stress the stability and robustness of online social behaviours. We will circle back to this research in the discussion section.
The idea that human social networks may be in equilibrium is relevant from the perspective of social behaviour because it suggests that, despite the complexity of individual interactions and the interplay between cognitive, emotional, and environmental factors, the overall structure of these networks remains stable over time. This stability points to a fundamental predictability in social systems, implying that, like physical systems, human networks follow certain rules, and their evolution is constrained. We will discuss how these constraints may arise from cognitive and material resource management, with individuals regulating their social ties to avoid overwhelming their social capacities. Besides, if a social network is in equilibrium, its properties can be predicted more accurately over time, facilitating intervention studies. Interventions can be introduced and compared against the equilibrium network, which serves as a baseline for analysis, helping to strengthen social cohesion or address social isolation. The possibility that this equilibrium behaviour may be general suggests that certain social behaviours and organization principles transcend cultural, situational or individual differences, connecting individual interactions and broader social structures. From a modelling perspective, equilibrium assumptions enable the use of simpler, more tractable mathematical models, such as statistical mechanics tools, developed for physical systems.
From a practical point of view, if social networks are in equilibrium, the potential for generalizability and robustness of findings from different studies significantly increases. Conclusions about the mechanisms driving network dynamics, structural trends, and other insights apply not only to a single observed point in time but to the entire unobserved evolution of the network after the transient formation period. This is specially relevant for methods that employ cross-sectional data to investigate network dynamics, like Exponential Random Graph Models. These approaches have an implicit and often overlooked assumption: the network must be in or near equilibrium, as discussed in [47]. In simple terms, when using cross-sectional data, the researcher assumes that every observation is statistically equivalent to any other, a condition that is fulfilled if the network is in equilibrium. This reduces the need for continuous data collection, lowering resource requirements and the burden on respondents.
Despite its importance, to the best of our knowledge, no prior research has explicitly focused on determining whether social networks are indeed in equilibrium. In this context, we present an example of an empirical social network that is in rigourous equilibrium, given by a rich dataset collected over four years. We also explore the implications of this finding, and, we discuss whether this should be a general phenomenon in any social network.
Figure 1: Representation of the network composition over the course of the 10 waves. The table in the upper part shows the presence of individuals across the waves. Each row represents a person, and each column represents a wave. A coloured cell indicates the person’s presence in that specific wave. The cell colour reflects the total number of waves that person participated in. Thus, all individuals present in n waves will have n cells coloured in the same colour, with a different colour for each n. colours are not evenly distributed because each individual is likely present in all waves of the same academic year (dropouts or new enrolments in the middle of the year are rare). On the right side, the percentages of people present in each number n of waves are displayed. For instance, 15.3% of individuals are present in 10 waves, 0.2% in 9 waves, etc. Below the table, a visual representation of the evolution of the network of positive and negative relationships is shown. Individuals are represented as nodes arranged in a circular layout, with blue links for positive relationships and orange links for negative ones. To illustrate the high turnover in network composition, all individuals are depicted in every wave, even if they have not yet joined or have already left the network.
2 Results
In this work, we analyse a dataset we have collected between 2020 and 2024 that contains the temporal evolution of the network of personal relationships among 888 people belonging to the Blas de Otero High School in Madrid (see the Materials and Methods section for details on the data collection, data composition and data curation). The reported relationships were coded as +2 (very good), +1 (good), -1 (bad), and -2 (very bad). Almost every student provided this list, resulting in the extraction of a weighted directed network for the entire high school, which we will refer to as snapshot or wave. We repeated this process every 16-20 weeks, and with this data we reconstructed the network at different points in time, collecting 10 snapshots of the network over 4 years. With respect to previous literature, we have introduced some improvements in the data collection process. Respondents are allowed to report an unlimited number of relationships, resulting in richer and more heterogeneous data. We also increased the sample size by one order of magnitude, with a larger number of snapshots, allowing for the detailed study of the dynamics in the long term while maintaining a certain level of granularity. Relationships are self-reported, and they are no longer binary. Instead, we include directions and weights to the relationships, between -2 and 2, allowing the consideration of negative relationships. Preliminary analyses of earlier, shorter versions of this dataset have been reported elsewhere [48, 41]. It is worth noting that by coding relationships into these four categories (-2, -1, +1 and +2), we inevitably lose many of the details that characterize social interactions. Thus, the resulting network we analyse in this study is not a direct reflection of the complexity of real personal relationships, but rather a simplified representation. Real-world relationships exist on a spectrum, with layers of emotional, social, and contextual subtleties that cannot be completely captured in these discrete categories. However, even with this simplification, we argue that identifying an equilibrium in this network serves as a meaningful proxy for the underlying dynamics of the real network. The equilibrium we detect in this reduced representation suggests that the macroscopic patterns of stability we observe likely reflect deeper robust social structures. If the simplified relationships follow this equilibrium dynamics, it implies that the more complex relationships they represent also tend toward stability over time.
In total, we surveyed 888 people, but not all of them were present in all waves because the composition of the network changes as time goes by. We surveyed people belonging to a high school over the course of four academic years. Every academic year, new people enter the high school to take the first course, and at the end of the school year, almost all the people belonging to the last course leave the high school (a minor portion of them needs to repeat the course if they fail their subjects). Thus, every academic year, some nodes appear in the network, and others disappear. In Figure 1 we depict this turnover in a visual manner. Interestingly, only 15% of the people are present both in the first and the last waves. This will become a very relevant fact when we show the network is in equilibrium, because this equilibrium will not result from the same group of people interacting over the course of four years, but rather an internal property of the network dynamics independent of the network composition. Due to this turnover, every snapshot of the network contains about 500 active nodes out of the 888.
Stationarity of the Transition Matrices
In a nutshell, the concept of dynamical equilibrium in a physical system implies that the macroscopic, average properties of the system remain constant, fluctuating around a stable value, while microscopic dynamics are actively changing. For instance, in a gas, the microscopic components (the particles) are constantly moving at different velocities, colliding, vibrating, etc. However, if the gas is in equilibrium, macroscopic properties such as temperature, pressure, or volume remain constant.
Drawing an analogy to a social system represented as a network, it is essential to define both the macroscopic properties and the microscopic dynamics. In a social network evolving over time, links appear, disappear, or change in nature. From the perspective of a node, an outgoing or incoming +1 link can become a +2, a -1, a -2 or disappear. We refer to an absent link as a 0 link. This constitutes the microscopic dynamics of the system: the evolution of individual ties/links. From these microscopic components, we can build macroscopic properties of the network. In the network, we define an edge as the connection between two nodes that contains two links, one from the first node to the second and one from the second node to the first. Since we have 5 types of links, -2, -1, 0, +1, +2, there are 25 possible edge types. Notice the distinction between link/tie and edge we introduce in our notation. Although it is usual to treat both as interchangeable, we keep this distinction throughout the paper. Our first macroscopic property is the distribution of different edge types: how many +2+2, +1+2, +1+1, etc., edges exist in the network. Notice that a +2+1 edge is not equivalent to a +1+2, specially from a dynamical perspective. Although both edges can evolve towards a +1-1 edge, in the +2+1 case the +2 needs to become a +1 and the +1 a -1, and in the +1+2 case, the +1 remains unchanged and the +2 becomes a -1. Therefore, we keep this distinction in all computations. A second macroscopic property is the in-degree and out-degree distributions of the nodes. For each type of link (+2, +1, -1, and -2), we count how many nodes have 0 incoming +2 links, 1 incoming +2 links, etc. This process is repeated for outgoing links and every type of link, generating eight degree distributions. A third macroscopic property can be the distribution of different triangle types that can form among three nodes. Other macroscopic properties include structural characteristics of the network, such as clustering, average shortest path length, assortativity, etc.
In this paper, we focus on the assessment of the degree, edge, and triangle distributions since we want to be able to define transition matrices to predict the expected equilibrium states of these macroscopic properties from the dynamics observed, granting the conclusions apply to all macroscopic properties. Before going into the details, it is important to comment on the limitations of our approach from the statistical mechanics point of view. The main challenge is the absence of a continuous concept of time in our analysis. By relying on snapshots taken every 20 weeks, we lose information about the transitions that occur between these intervals. Hence, there is an implicit assumption that no multiple transitions between edge states have occurred within the period between snapshots, minimizing the risk of hidden dynamics. It is clear that this assumption may not hold in all cases: In systems like a gas, for example, relevant changes happen on much shorter timescales, making such large gaps between observations inappropriate. However, in the context of social networks, the underlying dynamics evolve more slowly, and we believe that 20-week intervals provide a sufficient resolution. On the other hand, the use of statistical mechanical techniques in this study necessarily involves certain approximations to simplify the analysis of complex social systems. We aim at striking a balance between revealing meaningful sociological patterns and maintaining the rigour without becoming overly entangled in these mathematical details.
Let us introduce the concept of a transition matrix using the edges as our macroscopic property. In the case of edges, there are 25 possible edge states. If we take two consecutive snapshots of the network, for each individual edge, we can record what kind of transition it went through. A +2+2 edge can stay a +2+2, or become a +2+1, or even a +2-1, etc. If we repeat this process for all the links, we can construct a 25x25 matrix in which each row is the edge state before the transition, in the first snapshot, and each column is the edge state after the transition, in the second snapshot. The $ij$ element of this matrix would be the number of edges starting in state $i$ and ending in state $j$. In this matrix, we can divide each element by the total count in each row. By doing this, the $ij$ element in this final matrix represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state provided it started in the $i$ state. We call this matrix the transition matrix $P_{\beta\alpha}$ (also known as the Markov Matrix), from snapshot $\alpha$ to snapshot $\beta$. Additionally, we define the edge state distribution $\pi_{\alpha}$ as a column vector in which each element $i$ is the density of edges of type $i$ in the snapshot $\alpha$. With these two objects, it is direct to see that:
$\pi_{\beta}=P_{\beta\alpha}^{T}\pi_{\alpha}$ (1)
In other words, if we take the distribution of edges in one snapshot and multiply this distribution by the transition matrix, we obtain the distribution of edges in the next snapshot.
Since we have 10 snapshots, our data allows us to construct nine transition matrices between consecutive snapshots. The first relevant question is whether these nine transition matrices are statistically equivalent. If they were, it would indicate that the dynamics are stationary, meaning the transition probabilities between edge states remain constant over time. However, notice this does not necessarily imply that the system is in equilibrium. We could have stationary dynamics where macroscopic properties change over time, but these changes would be uniform because the transition probabilities do not vary.
To simplify the process, we check whether the nine transition matrices are statistically equivalent to the average transition matrix computed by averaging the nine individual matrices. To compare each transition matrix with the average, we resampled the individual matrix randomly from the data 1000 times. We calculated the distance between each element of the randomly sampled matrix and the corresponding element in the average transition matrix. To do such resampling, we created a pool with all the observed transitions for each individual edge in each pair of waves, removing the information about the actual waves in which that individual transition occurred. We then created each randomly resampled matrix by drawing samples at random from this pool with no information about the specific waves in which transitions occurred. We counted how often each element of the resampled matrix was farther from the corresponding element in the average matrix than the actual value for the matrix we wanted to compare. In summary, we calculated, element by element, the probability that a randomly sampled matrix element would be farther from the corresponding element of the average matrix than the same element in the actual transition matrix we were analysing.
[…]
It is worth noting that we perform an element-wise comparison. Although we could compare the matrices themselves using metrics like the Frobenius distance, we chose a granular approach to be able to identify individual differences. However, it is important to take into account that this method involves multiple comparisons of elements in a sampled matrix; specifically, we are comparing 625 elements against another 625 elements. Consequently, we expect some elements to deviate more than expected by chance, even if the matrices are equivalent. For such cases, we compute a z-score to determine how significant these deviations are from randomness.
[…]
Given these results, it is natural to question whether the network is changing significantly. One might doubt whether the observed stability in dynamics is due to the network barely changing, with most edges remaining constant. For this reason, we included panel $c)$ in Figure 2 to illustrate the main dynamics within the network. While some edges are indeed stable, such as the $+2+2$ edge, which is the most stable, there are still significant dynamics within these edges. Approximately half of the probability flow from this edge transitions to other states, showing that only a little over half of these edges persist from one wave to the next. For all other edges, the probability of transitioning to a different state is greater than that of remaining unchanged. This indicates that the dynamics are quite active, with a considerable turnover in edge states.
This strategy can also be applied to other macroscopic properties. For instance, at the level of triangles – similar to edges – we can identify all unique triangle states and construct a transition matrix for these states. Given that there are five different types of directed links, the number of combinations grows exponentially with the size of the macroscopic structures analysed. In the case of triangles, we identify 1924 unique triangles (the number we observe, not the theoretical maximum). Thus, we can construct a 1924x1924 transition matrix for triangles and perform a similar test.
[…]
Similarly, we can apply this method to degrees. For each node, we have eight types of degrees (in and out degrees for +2, +1, -1, and -2 links). For example, for the +2 in-degree, we examine how many nodes with an initial +2 in-degree transition to a different +2 in-degree in the next wave. This allows us to construct a transition matrix for +2 in-degrees. Although the matrix size varies depending on the specific degree analysed, the strategy remains consistent. We stress that, in all cases, we find that the transition matrices governing the evolution of these macroscopic properties are stationary, indicating that the probabilities driving the network’s evolution are constant over time.
Equilibrium State
As mentioned, the stationarity of the transition matrices does not necessarily imply that the network is in equilibrium; it may not even have an equilibrium state under these dynamics. As we will see, this is not our case.
Starting from equation 1, focusing on edges, we can determine whether the system has a stationary state under the dynamics governed by the average transition matrix calculated. If it exists, we can calculate and compare it with the actual state of the network. In equation 1, we obtain the edge abundances in the next wave by multiplying the transpose of the transition matrix by the edge abundances in the current wave. When the system reaches the stationary state, these edge abundances become invariant under the transpose of the transition matrix. In other words, to find the stationary state, we need to solve the following equation:
$\tilde{\pi}=P_{\beta\alpha}^{T}\tilde{\pi}$ (2)
Here, $\tilde{\pi}$ represents the stationary edge abundances, i.e., the edge abundances expected in the stationary state. Solving equation 2 is equivalent to finding the eigenvector of the transpose of the transition matrix associated with the eigenvalue 1. A comparable procedure can be applied to obtain the stationary abundances of triangles and the eight different degree distributions. Once we obtain these theoretical abundances of edges, triangles, and degrees expected in the stationary state under the system’s current dynamics, we can compare them with the current abundances to see where the system stands with respect to the stationary state. This comparison is depicted in Figure 3.
It is significant to note that in all cases, we find that the theoretical stationary states coincide with the observed empirical abundances, which indicates that the system has reached the stationary state. Nonetheless, to say that the system has reached the stationary state is not equivalent to say that the system has reached the equilibrium. To rigourously claim that the system is in dynamical equilibrium in the statistical mechanical sense, we need to check that the Detailed Balance Condition is fulfilled. The Detailed Balance Condition ensures that a system has reached equilibrium and is expressed as [49]:
$P(i)P(j|i)=P(j)P(i|j)$ (3)
Where $P(i)$ is the probability of the system being in state $i$, and $P(j|i)$ is the conditional probability of transitioning from state $i$ to state $j$ provided the system starts in state $i$. That is, this condition ensures that the probability flow for all transitions between every pair of states is equivalent in both directions. Using our notation:
$\tilde{\pi}_{i}(P_{\beta\alpha})_{ij}=\tilde{\pi}_{j}(P_{\beta\alpha})_{ji}$ (4)
Since both $\tilde{\pi}$ and $P_{\beta\alpha}$ are sampled from the data, there is some uncertainty associated with both sides of equation 4. Therefore, the Detailed Balance Condition needs to be fulfilled within the confidence intervals associated with both sides of the equation. Considering this, we find that equation 4 is satisfied for every transition between edge states. There is a small violation of the Detailed Balance Condition in five out of the 325 possible transitions (specifically, going from $+0+0$ to $+2+2$, $+1+2$, $+2+1$, $+1+1$, $+0+1$), but this violation is closely within the confidence intervals. Looking to the magnitude of the violation, we observe that it affects to only the 1.9%, 0.6% 0.4% 1.7% 3.6% of the transitions count involving these links, respectively. Thus, we consider that this violation barely disrupts the equilibrium of the network. It is possible to detect in Figure 3 the effect this violation has in the equilibrium abundances. For the edges and triangles affected, the equilibrium curves are slightly above the empirical abundances, although there is still an almost perfect overlap between both. The method’s ability to detect such a small violation supports the robustness of the equilibrium conditions found for the remaining transitions.
Discussion
In this paper, we have shown that the dynamical process driving the evolution of the personal relationships network of students belonging to a high school in Madrid, from 2020 to 2024, is stationary, and that the network itself is in dynamical equilibrium. We find that the abundances of edges, triangles and degrees are approximately the same in all snapshots and can be accurately predicted by the average transition matrix. Furthermore, a key feature of equilibrium in the statistical mechanics sense, Detailed Balance, is also verified with very small, point deviations.
We stress that this result goes far beyond what has been studied so far. As we mentioned in the introduction, some previous research reported a certain stability in the data analysed [20, 23, 17, 25, 29, 28, 13, 41], but none of them explored this stability further, and, to the best of our knowledge, no other work has shown rigorously that an empirical network is in equilibrium. In addition, as we mentioned in the introduction, other studies have found stability and robustness in social behaviour in broader populations. For instance, in [42], they show that social networks exhibit a resilience mechanism after the death of a friend, recovering the same number of active connections over time through increased interactions between friends of the decedent. This would be an example of how the network evolves to maintain its structural integrity. They also show that interactions between friends stabilize after a year following the death of a friend, similar to the stationarity in the transition probabilities in our findings. Moreover, they mention the possibility of the existence of a lower bound on individuals’ level of social connection, that would force them replace lost friendships more quickly than they are driven to establish friendships in general. Within the context of mobility patterns, in [43], the authors found that the number of locations an individual visits regularly is conserved over time, even while individual routines are unstable in the long term because of the continuous exploration of new locations. Furthermore, they find a connection between this number and the the Dunbar’s number, establishing a relation between these mobility patterns and the maintenance of their social relations. In [44, 45, 46] the authors explore how online social behaviour follows consistent patterns over time, with [45] focusing on the role played by ties of different strengths, and proving how different dynamical mechanisms compete for this stability to arise, an idea we will develop on the following discussion. Finally, in [50] the authors show how different group social behaviours are consistent within both students and adult groups, pointing towards the generality in the dynamical behaviour of social bonds.
While the specificities of our sample population may limit direct extrapolation to other social networks, we have reasons to believe that this result generalizes to broader populations. Apart from the previous studies mentioned, that report this robustness and stability of certain social behaviours, the exploration of the mechanisms behind this equilibrium dynamics can reveal fundamental principles of social network dynamics applicable to broader populations. The temporal evolution of any network of personal relationships is influenced mainly by three important factors, and we argue they are the ultimate responsibles for the arising of an equilibrium. The first factor is composed of endogenous mechanisms that control how and why these networks evolve at the level of links. As reviewed in [39], these mechanisms include a tendency towards the reciprocation of friendships, the creation of homophilic relationships (with both influence and selection mechanisms), the establishment of transitive and hierarchical structures, the formation of balanced triangles (triangles with an even number of negative links [51, 52]), the avoidance of conflicts, among others. Regarding the second factor, the top-level organizational structure of the society and the social foci shape the contexts in which relationships are formed, affecting who interacts with whom. In a high school, for instance, the fact that students share the same course, classes, or subjects biases the network structure. The third factor involves the cognitive constraints of individuals, and the social necessities they have. How people allocate their limited cognitive and material resources in their relationships to cover those necessities shapes the relational structure of each person, reflected in the creation of the typical Dunbar’s structures, as shown in [2, 3].
Within this theoretical framework, the concept of equilibrium arises naturally. Basically, in the evolution of personal relationships, the social foci, the context, and the organizational structure of the society conditions with whom we interact (and, hopefully, with whom we establish relationships). Once this context is stable, people establish relationships allocating their cognitive and material resources among the people with whom they share this social environment. In the broader picture of the network, dynamical mechanisms shape how these relationships are structured (in simple terms, a person allocates some resources to establish a relationship, and dynamical mechanisms determine who this person is within the broader network). Thus, every time a relationship disappears, there is a liberation of resources that can be reinvested in new relationships. This argument is supported by the findings we mentioned from [42]. At the level of dynamical mechanisms, there is a balancing process as well. For instance, when this new relationship is established, the reciprocation mechanisms will force it to be reciprocal. Nonetheless, there are other mechanisms destroying this reciprocity, like the formation of hierarchical structures. Similarly, there is a tendency towards the creation of balanced triangles, but also a tendency to avoid conflict and a high random component in conflict, that promotes the destruction of these triangles and the creation of unbalanced ones. Therefore, there is a competition between mechanisms having opposing effects on the network structures. The equilibrium is the expected outcome of the network dynamics.
An important insight from our results is the role the second factor plays in the observation of the equilibrium (the context, the social foci and the society’s organizational structures). Basically, to observe the network in equilibrium one needs to analyse the system at a proper temporal and spatial scale. The equilibrium would appear only after the network has had time to stabilize after the transient time, and the context of the observed network needs to be wide enough to be able to potentially saturate the social relationships of the people in the network. Let us illustrate this last point with an example. If we observe only a certain class within the school, it may happen that some people cease being friends within the class and compensate these relationships with people in other classes. However, if we are observing only this specific class, we may see how some relationships simply disappear from the system under our observation. If we draw an analogy with a physical system, to see the equilibrium it is necessary to observe the system within some ‘natural boundaries’ that contain all the relevant dynamics of the system. Thus, in a social system we reason that, to detect the dynamical equilibrium, it is necessary to observe the network within some ‘natural boundaries’, and these boundaries are defined by the context, social foci and organizational structure that is able to contain the majority of social relationships of the group. If not, you may observe certain stability in the system, but you may not be able to detect the equilibrium properly. In our case, the school seems to provide a perfect environment for this. Students may have other relationships outside the school, like familiar ties and contacts from extra-curricular activities, but the majority of their relationships are contained within the school, specially in these ages. Of course, the social foci can change in larger time scales. The organizational structure can slowly change, and in the case of a high school, at some point students leave, disrupting their network of relationships [53]. Nonetheless, within the high school, as a closed environment that saturates the possible relationships students can have, and observing the system at a proper temporal scale in which the context remains stable, this equilibrium exists as a consequence of the competition and balance between mechanisms, and between necessities and resources allocation.
We want to stress that in our dataset only 15% of nodes are present both in the first and the last waves. Thus, the equilibrium does not arise just because some fixed group of people coexist for a certain period of time. This result supports our view of the equilibrium as a consequence of the competition of dynamical mechanisms and a trade-off between resources allocation and social necessities more than a property of individuals themselves. Furthermore, this result indicates that statistically, people leaving and entering the school are similar, as the observed structure remains stable even when most of the network composition changes. This finding supports the view that there is not large variability in the structures of people’s individual relationships, as highlighted by [54], and our conclusions’ applicability to many other social networks.
If indeed social networks are generally in equilibrium, the findings of studies using cross-sectional data could potentially be generalized to the entire unobserved evolution of the network. When studied at the proper temporal and spatial scales after the transient, the observation of single points can provide robust information about the dynamical mechanisms driving the network evolution and the structural features observed. This provides solid ground for using methods like Exponential Random Graph Models and other methods that compare the network observation to null-models to understand the structural trends observed. From the social sciences perspective, this finding can ease the resource requirements to collect longitudinal data on social networks, reducing the burden on respondents. Also, it opens the door for the design of intervention studies. Since the network properties are stable over time, the researcher can introduce interventions and associate the observed changes to the intervention isolated from other potential drivers of the network evolution.
Finally, from a theoretical modelling perspective, this equilibrium has direct implications for the predictability of social behaviour and our understanding of the system’s interdependence. The resulting observed social networks are the product of competing mechanisms and resources management, and while relationships are dynamic and constantly evolving, influenced by cognitive, emotional and environmental factors, this evolution is highly constrained, and there is a tendency toward stable structures. The cognitive and emotional processes that govern social interactions lead to the formation of a robust social fabric. The development of simple mathematical models to describe human behaviour is justified by these insights. For instance, our result supports the application of statistical mechanical methods to social systems. Statistical mechanics often reveals universal patterns in physical systems, suggesting that human relationships may also exhibit this type of universal behaviours. This conclusion connects with the findings of [55], that deals with the predictability of societal changes, stressing the importance of combining such statistical models with domain-specific knowledge to make better predictions. Such an observation challenges the view that social systems are uniquely complex and unpredictable, raising the question of whether social phenomena can be explained by general laws or if they are inherently unique and context-dependent. To properly answer this question, future longitudinal studies should be conducted to establish whether this observed equilibrium is indeed a general property of social networks.
3 Materials and methods
In this section we provide details about the data collection process, the data composition and the data curation.
Data collection
The collection of our data was performed through surveys administered in the school via a computer interface. To elicit relationships, students were presented with a list of all other students in the high school. They were then asked to select individuals with whom they had a relationship. Specifically, the questionnaire included the following question: ‘You can now see the list of all the students in the school. Please mark those you have any relationship with by clicking ‘very good relationship’, ‘good relationship’, ‘bad relationship’ or ‘very bad relationship’. Only one choice is possible. If you do not mark any option, it will be understood to mean that you do not have a relationship with the person’. Typically, it took students about 15 minutes to complete the survey, and they were supervised by a school teacher throughout the process. Our Institutional Review Boards (IRB) stipulated an opt-out procedure. We should note that there were no opt-outs, effectively eliminating any potential selection bias. The only students who did not participate were those who were absent on the day of the experiment.
Data composition and data curation
In Table 1, we present the composition of the network snapshot by snapshot. The missing people column includes people who were absent on the day of the survey because there were no opt-outs. We removed some outliers from the analyses, defined as people with more than 30 outgoing very good relationships, more than 50 outgoing good relationships, more than 15 outgoing bad relationships or more than 15 outgoing very bad relationships. These numbers were selected by comparing outgoing with incoming degree distributions. The proportion of outliers removed is shown in the Outliers column. In any case, results with and without outliers are approximately equal, showing that their presence would not change our conclusions.
Table 1: Composition of each network snapshot. In the columns Sex, Missing Data and Outliers, numbers represent proportions.
Wave Respondents Sex (M/F) Missing Data Outliers
DEC 2020 409 0.52/0.48 0.11 0.03
MAY 2021 409 0.52/0.48 0.10 0.03
SEP 2021 530 0.49/0.51 0.06 0.04
FEB 2022 530 0.49/0.51 0.09 0.04
MAY 2022 530 0.49/0.51 0.00 0.07
SEP 2022 524 0.54/0.46 0.06 0.03
JAN 2023 535 0.54/0.46 0.09 0.07
MAY 2023 536 0.54/0.46 0.13 0.05
SEP 2023 554 0.51/0.49 0.06 0.04
JAN 2024 563 0.51/0.49 0.07 0.05
Data availability
All the aggregated data necessary to replicate our results can be found in this repository.
4 Acknowledgements
M.A.G.-C. acknowledges support from the Comunidad de Madrid through the grants for the hiring of pre-doctoral research personnel in training (reference PIPF-2023/COM-29487). M.A.G.-C. and A.S. acknowledge support from grant PID2022-141802NB-I00 (BASIC) funded by MCIN/AEI/10.13039/501100011033 and by ‘ERDFA way of making Europe’. A. S. T acknowledges support by FCT – Fundação para a Ciência e Tecnologia – through the LASIGE Research Unit, ref. UIDB/00408/2020 (https://doi.org/10.54499/UIDB/00408/2020) and ref. UIDP/00408/2020 (10.54499/UIDP/00408/2020)
5 Author Contributions
All authors conceived and conceptualized the research, A.S. collected the data, M.A.G.-C. curated the data, formalized the analyses and obtained the results, and all authors discussed and interpreted the results and wrote the manuscript.